# Bo Du1† Xuekang Zhu1,2† Xiaochen Ma3† Chenfan Qu4, 2† Kaiwen Feng1† Zhe Yang1 Chi-Man Pun5 Jian Liu2‡ Jizhe Zhou1‡

### v2

@@ 1 Introduction
However, the research efforts of FIDL have gradually split into four relatively independent research domains over time. 1) Deepfake detection (Deepfake) nguyen2019capsule ; li2018uadfv ; li2020xray ; dang2020ffd ; ni2022core ; cao2022recce ; yan2023ucf ; han2023fcd ; qian2020f3net ; liu2021spsl ; luo2021srm ; sun2022SIA ; zhuang2022uia ; shiohara2022sbi ; yan2024LSDA ; yan2023deepfakebench : detects human-centric manipulations such as face swapping, expression editing, or feature replacement. 2) Image manipulation detection/localization (IMDL) liu2020d ; MVSS_2021 ; trufor2023 ; liu2022pscc ; NCL_IML_2023 ; ma2023iml ; yu2024diffforensics ; zhu2025mesoscopic ; sparseViT_2025 : detects and localizes the tampering in natural images. 3) AI-Generated Image Detection (AIGC) univfd ; HiFi-Net2023 ; dualnet ; wang2023dire ; zhang2025towards : detects the images fully generated by deep generative models such as Stable Diffusion stablediffusion . 4) Document Image Manipulation Localization (Document)  psnet ; qu2023doctamper ; chen2024ffdn ; dong2024tifdm ; song2025caftb ; yu2023learning ; li2026document : localizes the forgery of various forms of document images, including receipts, certificates, and identification materials, with a particular focus on detecting modifications to the printed text.
@@
Deepfake Capsule-Net nguyen2019capsule (ICASSP19) VGG VGG_2015 Dynamic Routing Label Proposes a capsule network with dynamic routing and a VGG19 backbone.
@@
RECCE cao2022recce (CVPR22) Xception chollet2017xception Reconstruction Label Proposes a graph-based framework leveraging reconstruction differences
@@
SPSL liu2021spsl (CVPR21) Xception chollet2017xception Phase Spectrum Label Proposes phase-spectrum fusion with Xception for face forgery detection.
@@
UCF yan2023ucf (ICCV23) Xception chollet2017xception Multi-task Disentanglement Label Proposes multi-task disentanglement with Xception for deepfake generalization.
@@
SBI shiohara2022sbi (CVPR22) EfficientNet tan2019efficientnet Frequency,Blending Boundaries Label Proposes self-blended images to improve deepfake detection generalization.
@@
IMDL MVSS-Net Mantra_2019 (ICCV21) Resnet Resnet_2016 BayarConv,Sobel Label,Mask Exploit noise and boundary artifacts via multi-view learning for manipulation detection.
@@
CAT-Net CAT-Net2022 (IJCV22) HRNet HRNet DCT Mask Fuse RGB and DCT streams to learn compression artifacts for splice localization.
@@
PSCC-Net liu2022pscc (TCSVT22) HRNet HRNet Multi-Resolution Conv Label,Mask Progressively refine masks with spatio-channel correlations for high-reso localization.
@@
Trufor trufor2023 (CVPR23) Seformer SegFormer_2021 High-Reso,Multi-scale,Edge Label,Mask Fuse RGB and learned noise fingerprints to detect manipulations as anomalies.
@@
IML-ViT ma2023iml (Arxiv) ViT ViT_2021 BayarConv,SRM Filter Mask Use ViT with high-reso, multi-scale edge-aware design for manipulation localization.
@@
Mesorch zhu2025mesoscopic (AAAI25) Conv. Convnet_2022 ,Segfor. SegFormer_2021 DCT Mask Fuse micro and macro cues for mesoscopic image manipulation localization.
@@
AIGC Dire wang2023dire (ICCV23) Resnet Resnet_2016 Diffusion Reconstruction Label Use reconstruction error of diffusion for diffusion-generated images detection.
@@
DualNet dualnet (APSIPA23) CNN SRM,Low Frequency Label Fuse SRM residual and low-frequency content streams for AIGC detection.
@@
HiFiNet HiFi-Net2023 (CVPR23) HRNet HRNet Multi-branch Feature Extractor Label,Mask Learn hierarchical fine-grained representations of forgery attributes.
@@
Synthbuster bammey2023synthbuster (OJSP23) None Fourier Transform Label Leverage spectral artifacts in the frequency domain for diffusion detection.
@@
UnivFD univfd (CVPR23) CLIP-ViT clip None Label Use pretrained vision-language model features for unified detection.
@@
Document CAFTB song2025caftb (TOMM24) Resnet Resnet_2016 SRM Mask Proposes CAFTB-Net with dual-branch and cross-attention.
@@
TIFDM dong2024tifdm (TCE24) Resnet Resnet_2016 None Mask Proposes a robust network with multiscale attention.
@@
DTD qu2023doctamper (CVPR23) Conv. Convnet_2022 , Swin.Swin_2021 Frequency Mask Proposes DTD with frequency head and multi-view decoder.
@@
FFDN chen2024ffdn (ECCV24) ConvNext Convnet_2022 Wavelet, Frequency Mask Proposes FFDN combining visual enhancement and frequency decomposition
@@
Although individual benchmarks exist in some domains, such as DeepfakeBench yan2023deepfakebench for Deepfake and IMDLBenCo ma2024imdl for IMDL, a unified benchmark for all domains in FIDL remains blank. The absence of such a unified benchmark results in significant domain silos, where each domain independently constructs its datasets, models, and evaluation protocols without interoperability. Domain silos lead to redundant and uneven research across existing FIDL fields, and difficulty in establishing a general and unified FIDL approach, severely hindering the development of the entire FIDL field.
@@ 2 Related Works
DeepfakeBench yan2023deepfakebench is a Deepfake detection benchmark specifically designed to address the lack of uniformity in data processing pipelines, leading to inconsistent data inputs for detection models. IMDLBenCo ma2024imdl is a benchmark and codebase for IMDL, aiming to compare IMDL models through a unified training and evaluation protocol. AIGCDetectBenchmark EkkoAIGCDetectBenchmark is a repository for experiments on 9 AI-generated image detection methods.
@@ 3 ForensicHub
To enable broad interoperability and reduce duplication of effort, ForensicHub adopts an adapter-based design gamma1995design that ensures seamless integration with DeepfakeBench yan2023deepfakebench and IMDLBenCo ma2024imdl , two widely used benchmarks. This mechanism allows users to reuse existing models and datasets without major modification, while also supporting the definition of new models and benchmarks within ForensicHub under the unified protocol. This unified infrastructure simplifies cross-task benchmarking, supports reproducibility, and enables consistent evaluation across domains.
@@
Datasets used in this paper are: FaceForensics++ rossler2019faceforensics++ , Celeb-DF li2020celeb , DFD dolhansky2020dfd , FaceShifter li2019faceshifter and UADFV li2018uadfv for Deepfake; CASIA CASIA_2013 , COVERAGE Coverage_2016 , Columbia Columbia_2006 , IMD2020 IMD20_2020 , NIST16 NIST16_2019 , CocoGlide trufor2023 , and Autosplice jia2023autosplice for IMDL; DiffusionForensics wang2023dire , GenImage zhu2023genimage for AIGC; Doctamper qu2023doctamper , T-SROIE wang2022tsroie , OSTF qu2025ostf , TPIC-13 wang2022tpic , RTM luo2025rtm for Doc. A brief summary of each dataset is provided in Table 2, with more details in Appendix D.1.
@@
Deepfake FaceForensics++ rossler2019faceforensics++ 2019 Face 45,388 127,209 Label Automatic Train,Test
Celeb-DF-v1 li2020celeb 2020 Face 1,203 1,933 Label Automatic Train,Test
Celeb-DF-v2 li2020celeb 2020 Face 5,620 10,800 Label Automatic Train,Test
DeepFakeDetection dfd2019 2019 Face 10,741 91,800 Label Automatic Train,Test
DFDC dolhansky2020dfd 2020 Face 63,265 68,851 Label Automatic Train,Test
DFDCP dolhansky2019deepfake 2019 Face 5,901 11,321 Label Automatic Train,Test
FaceShifter li2020advancing 2020 Face 4,479 4,479 Label Automatic Train,Test
UADFV li2018uadfv 2018 Face 1,548 1,551 Label Automatic Train,Test
IMDL CASIAv1 CASIA_2013 2013 General 0 920 Label,Mask Manual Train
CASIAv2 CASIA_2013 2013 General 7,491 5,123 Label,Mask Manual Train
COVERAGE Coverage_2016 2016 General 100 100 Label,Mask Manual Train
Columbia Columbia_2006 2006 General 183 180 Label,Mask Automatic Train
IMD2020 CASIA_2013 2020 General 414 2,010 Label,Mask Manual Train
NIST16 NIST16_2019 2019 General 873 564 Label,Mask Manual Train
CocoGlide trufor2023 2023 General 512 512 Label,Mask Automatic Train
Autosplice jia2023autosplice 2023 General 2,273 3,621 Label,Mask Manual Train
AIGC DiffusionForensics wang2023dire 2023 General 134,000 481,200 Label Automatic Train,Val,Test
GenImage zhu2023genimage 2023 General 1,331,167 1,350,000 Label Automatic Train,Test
Document Doctamper qu2023doctamper 2023 Document 0 170,000 Mask Automatic Train,Test
OSTF qu2025ostf 2025 Document 16,405 3,685 Mask Manual Train,Test
RealTextManipulation luo2025rtm 2025 Document 52,314 17,423 Mask Automatic+Manual Train,Test
T-SROIE wang2022tsroie 2022 Document 21,268 4,326 Mask Manual Train,Test
Tampered-IC13 wang2022tpic 2022 Document 2,810 1,228 Mask Manual Train,Test
Models used in this paper are: Capsule-Net nguyen2019capsule , RECCE cao2022recce , SPSL liu2021spsl , UCF yan2023ucf , and SBI shiohara2022sbi for Deepfake detection; MVSS-Net MVSS_2021 , CAT-Net CAT-Net2022 , PSCC-Net liu2022pscc , Trufor trufor2023 , IML-ViT ma2023iml , and Mesorch zhu2025mesoscopic for image manipulation and localization; Dire wang2023dire , DualNet dualnet , HiFiNet HiFi-Net2023 , Synthbuster bammey2023synthbuster , and UnivFD univfd for AIGC detection; DTD qu2023doctamper , FFDN chen2024ffdn , CAFTB song2025caftb , TIFDM dong2024tifdm for document detection. These methods are from official repositories and our reimplementations. In addition, ForensicHub also selects 6 commonly used backbones in visual tasks, which are: Resnet Resnet_2016 , Xception chollet2017xception , EfficientNet tan2019efficientnet , Segformer SegFormer_2021 , Swin Transformer Swin_2021 , and ConvNext Convnet_2022 . Details about models can be found in Appendix D.2.
@@ 4 Benchmarks
In addition to being fully compatible with existing benchmarks, DeepfakeBench yan2023deepfakebench and IMDLBenCo ma2024imdl , ForensicHub further extends standardization efforts by introducing unified evaluation protocols for the AIGC and Document domains—two areas that previously lacked widely accepted benchmarks and codebases. We propose two protocols for two domains to evaluate generalization.
@@ 4.1 AI Generation Image Detection Benchmark
In the field of AIGC detection, the challenge in dataset construction usually lies not in obtaining a sufficient quantity of samples, since they can be easily generated using existing models, but in ensuring comprehensive coverage of a wide range of generative models. Therefore, we select only two commonly used public datasets: DiffusionForensics wang2023dire and GenImage zhu2023genimage . The former contains only diffusion-based generated images, while the latter covers a million-scale dataset constructed from eight SoTA generative models. Models are trained on DiffusionForensics and evaluated on different generative models within GenImage to assess generalization, as detection methods typically already achieve good performance on samples from the same generative model zhu2023genimage . The detailed data splits are summarized in Table D.1.3.
@@
ForensicHub implements five SoTA methods in AIGC detection: Dire wang2023dire , DualNet dualnet , HiFiNet HiFi-Net2023 , Synthbuster bammey2023synthbuster , and UnivFD univfd , among which Synthbuster has no official open-source code and is fully reimplemented by us. More details about models and training settings can be found in Appendix E.1.

### v3

# Bo Du1† Xuekang Zhu1,2† Xiaochen Ma3† Chenfan Qu4, 2† Kaiwen Feng1† Zhe Yang1 Chi-Man Pun5 Jian Liu2‡ Ji-Zhe Zhou1‡

@@ 1 Introduction
However, the research efforts of FIDL have gradually split into four relatively independent research domains over time. 1) Deepfake detection (Deepfake) Nguyen et al. (2019); Li and Lyu (2018); Li et al. (2020b); Dang et al. (2020); Ni et al. (2022); Cao et al. (2022); Yan et al. (2023a); Han et al. (2023); Qian et al. (2020); Liu et al. (2021a); Luo et al. (2021); Sun et al. (2022b); Zhuang et al. (2022); Shiohara and Yamasaki (2022); Yan et al. (2024b, 2023b): detects human-centric manipulations such as face swapping, expression editing, or feature replacement. 2) Image manipulation detection/localization (IMDL) Liu et al. (2020); Chen et al. (2021); Guillaro et al. (2023); Liu et al. (2022a); Zhou et al. (2023); Ma et al. (2023b); Yu et al. (2024); Zhu et al. (2025); Su et al. (2025): detects and localizes the tampering in natural images. 3) AI-Generated Image Detection (AIGC) Ojha et al. (2023); Guo et al. (2023); Xi et al. (2023); Wang et al. (2023); Zhang et al. (2025): detects the images fully generated by deep generative models such as Stable Diffusion Rombach et al. (2022). 4) Document Image Manipulation Localization (Document)  Shao et al. (2023); Qu et al. (2023); Chen et al. (2024); Dong et al. (2024); Song et al. (2025); Yu et al. (2023); Li et al. (2026): localizes the forgery of various forms of document images, including receipts, certificates, and identification materials, with a particular focus on detecting modifications to the printed text.
@@
Deepfake Capsule-Net Nguyen et al. (2019)(ICASSP19) VGG Simonyan and Zisserman (2015) Dynamic Routing Label Proposes a capsule network with dynamic routing and a VGG19 backbone.
@@
RECCE Cao et al. (2022)(CVPR22) Xception Chollet (2017) Reconstruction Label Proposes a graph-based framework leveraging reconstruction differences
@@
SPSL Liu et al. (2021a)(CVPR21) Xception Chollet (2017) Phase Spectrum Label Proposes phase-spectrum fusion with Xception for face forgery detection.
@@
UCF Yan et al. (2023a)(ICCV23) Xception Chollet (2017) Multi-task Disentanglement Label Proposes multi-task disentanglement with Xception for deepfake generalization.
@@
SBI Shiohara and Yamasaki (2022)(CVPR22) EfficientNet Tan and Le (2019) Frequency,Blending Boundaries Label Proposes self-blended images to improve deepfake detection generalization.
@@
IMDL MVSS-Net Wu and others (2019)(ICCV21) Resnet He et al. (2016) BayarConv,Sobel Label,Mask Exploit noise and boundary artifacts via multi-view learning for manipulation detection.
@@
CAT-Net Kwon et al. (2022)(IJCV22) HRNet Wang et al. (2020) DCT Mask Fuse RGB and DCT streams to learn compression artifacts for splice localization.
@@
PSCC-Net Liu et al. (2022a)(TCSVT22) HRNet Wang et al. (2020) Multi-Resolution Conv Label,Mask Progressively refine masks with spatio-channel correlations for high-reso localization.
@@
Trufor Guillaro et al. (2023)(CVPR23) Seformer Xie et al. (2021) High-Reso,Multi-scale,Edge Label,Mask Fuse RGB and learned noise fingerprints to detect manipulations as anomalies.
@@
IML-ViT Ma et al. (2023b)(Arxiv) ViT Dosovitskiy et al. (2021) None Mask Use ViT with high-reso, multi-scale edge-aware design for manipulation localization.
@@
Mesorch Zhu et al. (2025)(AAAI25) Conv. Liu et al. (2022b),Segfor. Xie et al. (2021) DCT Mask Fuse micro and macro cues for mesoscopic image manipulation localization.
@@
AIGC Dire Wang et al. (2023)(ICCV23) Resnet He et al. (2016) Diffusion Reconstruction Label Use reconstruction error of diffusion for diffusion-generated images detection.
@@
DualNet Xi et al. (2023)(APSIPA23) CNN SRM,Low Frequency Label Fuse SRM residual and low-frequency content streams for AIGC detection.
@@
HiFiNet Guo et al. (2023)(CVPR23) HRNet Wang et al. (2020) Multi-branch Feature Extractor Label,Mask Learn hierarchical fine-grained representations of forgery attributes.
@@
Synthbuster Bammey (2023)(OJSP23) None Fourier Transform Label Leverage spectral artifacts in the frequency domain for diffusion detection.
@@
UnivFD Ojha et al. (2023)(CVPR23) CLIP-ViT Radford et al. (2021) None Label Use pretrained vision-language model features for unified detection.
@@
Document CAFTB Song et al. (2025)(TOMM24) Resnet He et al. (2016) SRM Mask Proposes CAFTB-Net with dual-branch and cross-attention.
@@
TIFDM Dong et al. (2024)(TCE24) Resnet He et al. (2016) None Mask Proposes a robust network with multiscale attention.
@@
DTD Qu et al. (2023)(CVPR23) Conv. Liu et al. (2022b), Swin.Liu et al. (2021b) Frequency Mask Proposes DTD with frequency head and multi-view decoder.
@@
FFDN Chen et al. (2024)(ECCV24) ConvNext Liu et al. (2022b) Wavelet, Frequency Mask Proposes FFDN combining visual enhancement and frequency decomposition
@@
Although individual benchmarks exist in some domains, such as DeepfakeBench Yan et al. (2023b) for Deepfake and IMDLBenCo Ma et al. (2024) for IMDL, a unified benchmark for all domains in FIDL remains blank. The absence of such a unified benchmark results in significant domain silos, where each domain independently constructs its datasets, models, and evaluation protocols without interoperability. Domain silos lead to redundant and uneven research across existing FIDL fields, and difficulty in establishing a general and unified FIDL approach, severely hindering the development of the entire FIDL field.
@@ 2 Related Works
DeepfakeBench Yan et al. (2023b) is a Deepfake detection benchmark specifically designed to address the lack of uniformity in data processing pipelines, leading to inconsistent data inputs for detection models. IMDLBenCo Ma et al. (2024) is a benchmark and codebase for IMDL, aiming to compare IMDL models through a unified training and evaluation protocol. AIGCDetectBenchmark Zhong et al. (2023) is a repository for experiments on 9 AI-generated image detection methods.
@@ 3 ForensicHub
To enable broad interoperability and reduce duplication of effort, ForensicHub adopts an adapter-based design Gamma (1995) that ensures seamless integration with DeepfakeBench Yan et al. (2023b) and IMDLBenCo Ma et al. (2024), two widely used benchmarks. This mechanism allows users to reuse existing models and datasets without major modification, while also supporting the definition of new models and benchmarks within ForensicHub under the unified protocol. This unified infrastructure simplifies cross-task benchmarking, supports reproducibility, and enables consistent evaluation across domains.
@@
Datasets used in this paper are: FaceForensics++ Rossler et al. (2019), Celeb-DF Li et al. (2020c), DFD Dolhansky et al. (2020), FaceShifter Li et al. (2019) and UADFV Li and Lyu (2018) for Deepfake; CASIA Dong et al. (2013), COVERAGE Wen et al. (2016), Columbia Hsu and Chang (2006), IMD2020 Novozamsky et al. (2020), NIST16 Guan et al. (2019), CocoGlide Guillaro et al. (2023), and Autosplice Jia et al. (2023) for IMDL; DiffusionForensics Wang et al. (2023), GenImage Zhu et al. (2023) for AIGC; Doctamper Qu et al. (2023), T-SROIE Wang et al. (2022b), OSTF Qu et al. (2025), TPIC-13 Wang et al. (2022a), RTM Luo et al. (2025) for Doc. A brief summary of each dataset is provided in Table 2, with more details in Appendix D.1.
@@
Deepfake FaceForensics++ Rossler et al. (2019) 2019 Face 45,388 127,209 Label Automatic Train,Test
Celeb-DF-v1 Li et al. (2020c) 2020 Face 1,203 1,933 Label Automatic Train,Test
Celeb-DF-v2 Li et al. (2020c) 2020 Face 5,620 10,800 Label Automatic Train,Test
DeepFakeDetection Google AI Blog (2019) 2019 Face 10,741 91,800 Label Automatic Train,Test
DFDC Dolhansky et al. (2020) 2020 Face 63,265 68,851 Label Automatic Train,Test
DFDCP Dolhansky et al. (2019) 2019 Face 5,901 11,321 Label Automatic Train,Test
FaceShifter Li et al. (2020a) 2020 Face 4,479 4,479 Label Automatic Train,Test
UADFV Li and Lyu (2018) 2018 Face 1,548 1,551 Label Automatic Train,Test
IMDL CASIAv1 Dong et al. (2013) 2013 General 0 920 Label,Mask Manual Train
CASIAv2 Dong et al. (2013) 2013 General 7,491 5,123 Label,Mask Manual Train
COVERAGE Wen et al. (2016) 2016 General 100 100 Label,Mask Manual Train
Columbia Hsu and Chang (2006) 2006 General 183 180 Label,Mask Automatic Train
IMD2020 Dong et al. (2013) 2020 General 414 2,010 Label,Mask Manual Train
NIST16 Guan et al. (2019) 2019 General 873 564 Label,Mask Manual Train
CocoGlide Guillaro et al. (2023) 2023 General 512 512 Label,Mask Automatic Train
Autosplice Jia et al. (2023) 2023 General 2,273 3,621 Label,Mask Manual Train
AIGC DiffusionForensics Wang et al. (2023) 2023 General 134,000 481,200 Label Automatic Train,Val,Test
GenImage Zhu et al. (2023) 2023 General 1,331,167 1,350,000 Label Automatic Train,Test
Document Doctamper Qu et al. (2023) 2023 Document 0 170,000 Mask Automatic Train,Test
OSTF Qu et al. (2025) 2025 Document 16,405 3,685 Mask Manual Train,Test
RealTextManipulation Luo et al. (2025) 2025 Document 52,314 17,423 Mask Automatic+Manual Train,Test
T-SROIE Wang et al. (2022b) 2022 Document 21,268 4,326 Mask Manual Train,Test
Tampered-IC13 Wang et al. (2022a) 2022 Document 2,810 1,228 Mask Manual Train,Test
Models used in this paper are: Capsule-Net Nguyen et al. (2019), RECCE Cao et al. (2022), SPSL Liu et al. (2021a), UCF Yan et al. (2023a), and SBI Shiohara and Yamasaki (2022) for Deepfake detection; MVSS-Net Chen et al. (2021), CAT-Net Kwon et al. (2022), PSCC-Net Liu et al. (2022a), Trufor Guillaro et al. (2023), IML-ViT Ma et al. (2023b), and Mesorch Zhu et al. (2025) for image manipulation and localization; Dire Wang et al. (2023), DualNet Xi et al. (2023), HiFiNet Guo et al. (2023), Synthbuster Bammey (2023), and UnivFD Ojha et al. (2023) for AIGC detection; DTD Qu et al. (2023), FFDN Chen et al. (2024), CAFTB Song et al. (2025), TIFDM Dong et al. (2024) for document detection. These methods are from official repositories and our reimplementations. In addition, ForensicHub also selects 6 commonly used backbones in visual tasks, which are: Resnet He et al. (2016), Xception Chollet (2017), EfficientNet Tan and Le (2019), Segformer Xie et al. (2021), Swin Transformer Liu et al. (2021b), and ConvNext Liu et al. (2022b). Details about models can be found in Appendix D.2.
@@ 4 Benchmarks
In addition to being fully compatible with existing benchmarks, DeepfakeBench Yan et al. (2023b) and IMDLBenCo Ma et al. (2024), ForensicHub further extends standardization efforts by introducing unified evaluation protocols for the AIGC and Document domains—two areas that previously lacked widely accepted benchmarks and codebases. We propose two protocols for two domains to evaluate generalization.
@@ 4.1 AI Generation Image Detection Benchmark
In the field of AIGC detection, the challenge in dataset construction usually lies not in obtaining a sufficient quantity of samples, since they can be easily generated using existing models, but in ensuring comprehensive coverage of a wide range of generative models. Therefore, we select only two commonly used public datasets: DiffusionForensics Wang et al. (2023) and GenImage Zhu et al. (2023). The former contains only diffusion-based generated images, while the latter covers a million-scale dataset constructed from eight SoTA generative models. Models are trained on DiffusionForensics and evaluated on different generative models within GenImage to assess generalization, as detection methods typically already achieve good performance on samples from the same generative model Zhu et al. (2023). The detailed data splits are summarized in Table D.1.3.
@@
ForensicHub implements five SoTA methods in AIGC detection: Dire Wang et al. (2023), DualNet Xi et al. (2023), HiFiNet Guo et al. (2023), Synthbuster Bammey (2023), and UnivFD Ojha et al. (2023), among which Synthbuster has no official open-source code and is fully reimplemented by us. More details about models and training settings can be found in Appendix E.1.